OpenStudy (anonymous):

|7-2q|=3 there are 2 answers. please show work.

OpenStudy (anonymous):

|7-2q|=3 is 7-2q=3 or 7+2q=3 -2q=-4 or 2q=-4 q = 2 or q = -2